OpenStudy (anonymous):

How would you factor 2x3 – 24x2y + 12xy – 144y2 using the grouping method?

OpenStudy (anonymous):

\[2x^3 - 24x^2y + 12xy - 144y^2 = 2x^2(x -12y) +12y(x-12y)\] \[ = (x-12y)(2x^3 + 12y)\]

OpenStudy (anonymous):

Can you write out how you did this, so I can understand a little bit better?

OpenStudy (anonymous):

See what you get if you take common from: \[\huge 2x^3 - 24x^2y\]